P(X = x) = {

1

3

x, if x ∈ {1,2}

0, otherwise.


1
Expert's answer
2021-02-24T06:09:54-0500

Given probability function is-

"P(X=x)=3x+1,x\in" {"1,2" }

"=0,\text{otherwise}"


At "x=0,P(1)=1"


At "(x=1),P(1)=3(1)+1=4"

At "(x=2),P(2)=3(2)+1=7"


Hence the Probability Distribution is given by-